OpenStudy (gabylovesyou):

Which of the following is a solution of 4x^2 = - 9x - 4?

OpenStudy (gabylovesyou):

9 plus or minus the square root of 17 divided by 8. negative 9 plus or minus the square root of 145 divided by 8. 9 plus or minus the square root of 145 divided by 8. negative 9 plus or minus the square root of 17 divided by 8

OpenStudy (anonymous):

x =(-9-√17)/8=-1.640 x =(-9+√17)/8=-0.610

undeadknight26 (undeadknight26):

Wolframalpha.com is better ;)

OpenStudy (anonymous):

ill check it out thx

OpenStudy (just_one_last_goodbye):

need help? ^_^

OpenStudy (gabylovesyou):

yes i do need help :)

OpenStudy (just_one_last_goodbye):

is that a negative 9 or subtract 9?

OpenStudy (anonymous):

like i just gave you 2 possible solutions!! xD

OpenStudy (gabylovesyou):

negative 9

undeadknight26 (undeadknight26):

#DeadPool117 Ftw He gave a direct answer but still...

OpenStudy (anonymous):

yea :3

undeadknight26 (undeadknight26):

Try explaining though :D I might have to write you up ;)

OpenStudy (just_one_last_goodbye):

ready for the steps @Gabylovesyou

OpenStudy (gabylovesyou):

@Deadpool117 yes please explain

OpenStudy (anonymous):

Ok ill explain...

OpenStudy (gabylovesyou):

yes @just_one_last_goodbye

OpenStudy (anonymous):

1. Simplify -4x2-9x - 4

OpenStudy (just_one_last_goodbye):

#1 subtract (-9x -4) from both sides

OpenStudy (anonymous):

-4x2 - 9x - 4 = -1 • (4x2 + 9x + 4)

OpenStudy (just_one_last_goodbye):

The second step can be followed in 2 way either completing the square or Quadratic Formula.

OpenStudy (anonymous):

let her explain, its kindof hard to...

OpenStudy (just_one_last_goodbye):

@Gabylovesyou please choose ^_^

OpenStudy (just_one_last_goodbye):

@Deadpool117 im a guy o-o

OpenStudy (anonymous):

O_O.......

OpenStudy (just_one_last_goodbye):

"My little brony all dressed in sexy" c;

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

I lost 11 fans :O

OpenStudy (just_one_last_goodbye):

@Gabylovesyou let just do squaring ^_^

OpenStudy (just_one_last_goodbye):

after the first step this is what our problem should look like now\[4x^2 + 9x +4 =0\]

OpenStudy (gabylovesyou):

ok for step 2, what is easier ?

OpenStudy (just_one_last_goodbye):

well squaring is easier ^_^

OpenStudy (gabylovesyou):

ok then squaring :)

OpenStudy (just_one_last_goodbye):

Ok next step is dividing both sides by 4

OpenStudy (gabylovesyou):

how do i do that ..

OpenStudy (gabylovesyou):

@just_one_last_goodbye

OpenStudy (just_one_last_goodbye):

sorry i was in a Live Lesson

OpenStudy (gabylovesyou):

@just_one_last_goodbye Don't worry :)

OpenStudy (just_one_last_goodbye):

Do you know where I got the 4 from or no?

OpenStudy (gabylovesyou):

cuz of 4x^2 ?

OpenStudy (just_one_last_goodbye):

yes correct ^_^

OpenStudy (just_one_last_goodbye):

so when you divide both sides by 4 you get\[x^2 + \frac{ 9x }{ 4 } +1 = 0\]

OpenStudy (gabylovesyou):

ok.. :)

OpenStudy (just_one_last_goodbye):

Then subtract 1 from both sides \[x^2 +9\ \frac{ 9x }{ 4 } =-1\]

OpenStudy (gabylovesyou):

ok

OpenStudy (just_one_last_goodbye):

._. seems like OS doesn't have this but its a triple fraction

OpenStudy (just_one_last_goodbye):

Add \[(9/4/2)^2\] to both sides

OpenStudy (just_one_last_goodbye):

^after that you'll get this\[x^2 +\frac{ 9x }{ 4 } +\frac{ 81 }{ 64 } = \frac{ 17 }{ 64 }\]

OpenStudy (gabylovesyou):

ok

OpenStudy (anonymous):

@Gabylovesyou @just_one_last_goodbye ya'll are cute

undeadknight26 (undeadknight26):

I am cute too D: Don't forget me ;-;

OpenStudy (anonymous):

@undeadknight26 yes you are

OpenStudy (gabylovesyou):

@jeremydaigle101 Thanks hun ^-^ @undeadknight26 Why yes, yes u r ;p

OpenStudy (anonymous):

@Gabylovesyou you are welcome

undeadknight26 (undeadknight26):

XD <3 Your cuter ;)

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

how old are ya'll i am 15

OpenStudy (anonymous):

What is the solution of this system of linear equations? http://media.education2020.com/evresources/2003-11-01-00-00_files/i0230000.jpg A. (1, 0) B. (0, 0) C. (0, 1) D. x = 0

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

can you help?